### v1

@@ 5 Metallicity properties
We determined metallicities ([Fe/H]) for the individual stars from the equivalent widths of the near-IR Ca II triplet (CaT) lines. Since the available calibrations are developed for RGB stars, we will be restricting our analysis to this sub-sample (i.e., , with , which helps separating them from the hotter evolving giants). We verified the goodness of this selection using a set of isochrones (Bressan et al. 2012) with ([Fe/H] dex) and Gyr, which traced the lower range of metallicities and stellar ages for the RGB stars as obtained from the SFH analysis of IC 1613 (Skillman et al. 2014). We also selected all stars with S/N to keep the uncertainties on the equivalent widths (EWs) less than 20%. This resulted in a sub-sample of 275 input sources.
@@ 6 Summary and future work
The quality of the dataset allowed for a generally accurate spectral classification (i.e., with determined to better than 500 K) of 808 stars. We found a majority of K-type stars, and representatives from all the other types (from O- to M-). The classification included the identification of a sample of 24 probable Be stars, together with a sample of 14 probable C stars.
@@
For all types except the Be stars, the l.o.s. velocities have been determined with a pixel-by-pixel analysis, resulting in measurements with a mean precision of km and S/N For the Be stars we have determined the velocities by cross-correlation with dedicated templates. From this sample we identified 746 probable members (i.e., having ), which show a significant rotation pattern (when modelled as a linear velocity gradient, this yields km s arcmin). This is the first time that rotation for the stellar component is detected with high significance for this galaxy.
@@
We further analysed the probable members looking at the kinematic properties of each individual field and for different stellar populations. The individual fields analysis confirmed the presence of a rotation signal, and the decrease with radius of the velocity dispersion values that appears to flatten moving outward. We found a general agreement with the velocity field of the neutral HI component, as analysed in Read et al. (2016), though, on average, the stellar values showed higher velocity dispersion and a slower rotation, which can be ascribed to asymmetric drift. The kinematic analysis of the different selected stellar populations showed a coherent picture with the analysis of the main sample, with the MS stars following more closely the kinematics of the HI gas than the evolved red giant stars. It appears that the kinematics of this younger stellar component is still coupled to that of the gas.
@@
The chemical analysis was conducted on the sub-sample of RGB stars exploiting the wavelength region of the Ca II triplet lines. We obtained a median dex, with scatter values of dex and dex, in fairly good agreement with previous results reported by Kirby et al. (2013). The average [Fe/H] is well within the rms scatter of the stellar luminosity-metallicity relation found for the LG dwarf galaxies by the same authors. The spatial distribution of the [Fe/H] values showed no significant signs of a radial metallicity gradient, compatible with results from other LG dwarf galaxies (Leaman et al. 2013; Taibi et al. 2022).
@@
In future dedicated efforts, we plan to perform a dedicated analysis of the massive stars in our sample using the FASTWIND stellar atmosphere code (Puls et al. 2005), which will allow us to obtain reliable stellar parameters from their spectra. We also aim to improve the identification process of the M-type stars in our sample, and therefore to constrain the spectroscopic C/M ratio. This will tell us more of the chemical properties of the intermediate-age stellar population of IC 1613. Finally, this MUSE data-set could also be used to analyse the ionized gas medium using available emission lines, like or nebular oxygen lines in order to get an insight on the kinematics of the shell-like gas structure visible in field F1 and compare it with that of the youngest stars.
@@ Acknowledgements.
We thank M.-R. Cioni, L. R. Patrick and J. Read for useful discussions and comments. ST acknowledges funding of a Leibniz-Junior Research Group (PI: M. Pawlowski; project number J94/2020) via the Leibniz Competition. GB and CG acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia en Innovación (AEI-MICIN) and the European Regional Development Fund (ERDF) under Grant Number PID2020-118778GB-I00/10.13039/50110001103 and the AEI under Grant Number CEX2019-000920-S. GI acknowledges financial support from the European Research Council (ERC) through the ERC Consolidator Grant DEMOBLACK, under contract No. 770017. SK acknowledges funding from UKRI in the form of a Future Leaders Fellowship (grant no. MR/T022868/1) This research has made use of NASA’s Astrophysics Data System, VizieR catalogue access tool (CDS, Strasbourg, France, DOI: 10.26093/cds/vizier), and extensive use of Python3.8 (Van Rossum & Drake 2009), including iPython (v8.3, Pérez & Granger 2007), Numpy (v1.21, Harris et al. 2020), Scipy (v1.7, Virtanen et al. 2020), Matplotlib (v3.5, Hunter 2007), Astropy (v5.0, Astropy Collaboration et al. 2022) and Scikit-learn (v1.0, Pedregosa et al. 2011) packages.References
@@
- Demers et al. (2006) Demers, S., Battinelli, P., & Kunkel, W. E. 2006, ApJ, 636, L85
@@
- Swan et al. (2016) Swan, J., Cole, A. A., Tolstoy, E., & Irwin, M. J. 2016, MNRAS, 456, 4315

### v2

@@ 5 Metallicity properties
The chemical analysis focuses only on the RGB stars in our sample, because, due to the low resolution of the MUSE data, it is difficult to make a quantitative estimate of spectral parameters other than using the spectral fitting techniques implemented in this work (see details in Appendix B.1). Therefore, we determined metallicities ([Fe/H]) for individual RGB stars from the equivalent widths of the near-IR Ca II triplet (CaT) lines. We restrict our analysis to this subsample by selecting stars with and , which helps separating them from the hotter evolving giants. We verified the goodness of this selection using a set of isochrones (Bressan et al. 2012) with ([Fe/H] dex) and Gyr, which traced the lower range of metallicities and stellar ages for the RGB stars as obtained from the SFH analysis of IC 1613 (Skillman et al. 2014). We also selected all stars with S/N to keep the uncertainties on the equivalent widths (EWs) less than 20%. This resulted in a sub-sample of 275 input sources.
@@
6 Discussion
In the previous sections, we have provided the results of the kinematic and chemical analysis of the MUSE data of IC 1613. The general picture we get is that both the gaseous and stellar components are rotating, with the evolved stars appearing to decouple from the motion of the HI mainly due to the asymmetric drift. The MS stars, on the other hand, appear to have maintained kinematics similar to that of the neutral gas. The velocity dispersion profile decreases with radius for all tracers, although the central MS value may be affected by the presence of binaries or the expansion motion of the central hole, which may bias it towards a higher value. The lack of neutral gas in the central parts is also the probable cause of the sudden decrease in the inner HI velocity dispersion. Nevertheless, there seems to be a correlation between the velocity dispersion of the different tracers and their age, with values increasing as we move from the gas towards the evolved stars at all radii. At the same time, the evolved stars seem to have a lower rotation support than the HI and MS stars.
In the following, we examine these age-kinematic relations in detail, discussing their implications for the formation mechanisms of IC 1613. At the same time, we address the fact that the inclination of this galaxy is rather uncertain, proposing an alternative way of calculating it using only stellar tracers of different ages. Although we do not obtain conclusive results, we find the multi-tracer analysis to be a promising way to resolve this issue.
6.1 Age-kinematics relations
In Sect. 4.2 and 4.3, we have shown how the different stellar tracers inspected exhibit different kinematic behaviour. Here we make more explicit the connection of these changes with stellar age, also taking into account information recovered from the chemical analysis. In particular, as shown in Fig. 10, we inspect the variation of the velocity dispersion and of the rotation-to-dispersion support ratio as a function of age.
In the first three panels of Fig. 10 (from left to right, from top to bottom), we show the age-velocity dispersion variation of the different kinematic tracers. We note that each point has an age value that is only the average of the probable age range covered by the individual tracers. To provide a finer sampling of the age axis, we added two additional subsamples selected from the RGB stars with metallicity measurements. We used their median value (see Table 5) to divide them into a metal-poor (MP) and a metal-rich (MR) subsample. According to the age-metallicity relation obtained from the SFH of IC 1613 (Skillman et al. 2014), they cover an age range between and , respectively. In all three panels, the values of each tracer are reported in Table 4, while for the HI we took the average of over the entire radial range (top left panel), the average for arcmin (top right panel) and the average for arcmin (bottom left panel). The average of the profile measured for other galaxies in the literature is also shown for comparison.
From the figure, it can be seen that the velocity dispersion tends to increase as we move towards higher ages, both when considering each tracer in its entirety as well as when considering the central and external pointings separately. In particular, we see that in the top two panels the stellar is higher than , but increases only slightly towards older ages, being consistent with a constant trend when considering the errors. In the bottom-left panel, however, we see a clear linear increase, with the of the younger tracers being significantly lower than . Considering the presence of the central HI hole, whose causes may have affected not only the kinematics of the HI itself but also that of the MS stars, the external pointings probably provide a more accurate picture of the age-velocity dispersion relation of IC 1613. In the bottom-right panel of Fig. 10, we show instead the ratio of rotational to dispersion support (without accounting for the inclination) for both stars and gas as a function of age. The HI value is calculated at the effective radius , while the stellar values are the average of the external F2 and F3 pointings (see again Table 4). Again, a linear trend is evident, with decreasing with age, with values between .
Such age-kinematics trends are in good agreement with those of two similarly bright, gas-rich and isolated LG dwarf galaxies, WLM and NGC 6822. Both systems show a that decreases with age and a that grows with age (Leaman et al. 2012; Belland et al. 2020)181818Belland et al. (2020) found such trends in [Fe/H], which also in the case of NGC 6822 are a good proxy for age.. In particular, an age-velocity dispersion relation (AVR) has also been observed at higher luminosities for the LMC, M33, M31 and the MW, while the satellite dSphs of the MW seem to show a constant trend with age (e.g. Leaman et al. 2017, but also Dorman et al. 2015; Quirk et al. 2019, 2022; and e.g. Poci et al. 2019; Shetty et al. 2020; Pessa et al. 2023, for AVR studies outside the LG). The AVR has historically been interpreted by assuming that stars are born from a gas that is always in a dynamically cold state, with the increase in velocity dispersion with age due to various scattering mechanisms (e.g. giant molecular clouds or spiral arms) that secularly increase the orbital energy (Spitzer & Schwarzschild 1951, 1953; Carlberg et al. 1985; Kokubo & Ida 1992). Alternatively, the AVR could be interpreted by assuming that stars are born from an interstellar medium that becomes less turbulent over time due to a decreased gas fraction and less gravitational instability in the gas disc (e.g. Bournaud et al. 2009; Forbes et al. 2012). The analysis of Leaman et al. (2017) favours this latter scenario, in which a latent dynamical heating also seems to be at play moving towards higher galactic masses, while environmental effects may have affected the MW satellite systems. Although the coarse age resolution of our sample prevents us from performing a more quantitative analysis of the AVR of IC 1613, our results seem to support the conclusions of Leaman et al. (2017).
A similar argument could be made inspecting the variation of the rotation support with age. We have seen that the of IC 1613 decrease with age, similarly to WLM and NGC 6822. However, few other systems in the LG have kinematic studies covering the youngest stars, with the majority focusing on the more accessible red giant population. In this context, Wheeler et al. (2017) showed with an homogeneous kinematic analysis that in general the gas-rich systems of the LG are not strong rotators in their evolved stellar component (i.e. they have ). These results appear valid regardless of the environment, in agreement with recent results found for dwarf galaxies residing in voids (de los Reyes et al. 2023). The average rotation-to-dispersion support of the evolved stars in IC 1613 is , which remains even taking into account the uncertainties over the inclination value (see Sect. 6.2). This is in agreement with the results of Wheeler et al. (2017) who concluded, comparing with zoom-in cosmological simulations of isolated dwarf galaxies, that dwarf galaxies probably form as puffy dispersion-supported systems, rather than with a cold kinematics. Considering that IC 1613 evolved mostly in isolation, the age trends observed in and appear to be intrinsic features of this system, the interpretation of which favours the Leaman et al. (2017) and Wheeler et al. (2017) scenarios.
The age-kinematic trends of IC 1613 are also in broad agreement with observations of galaxies at different redshift (), where it can be seen that in general increases, while decreases, as a function of (e.g. Wisnioski et al. 2015, 2019; Übler et al. 2019). These studies, however, barely cover the mass regime of dwarf galaxies, while recent results seem to question whether high redshift galaxies are dispersion-supported, with a similar to that of local spirals (Rizzo et al. 2020, 2021; Fraternali et al. 2021; Lelli et al. 2021). Such systems are more massive than local dwarf galaxies, and at present only a few facilities such as the James Webb space telescope, can help to understand if rotation-supported dwarfs were common in the early Universe (e.g. de Graaff et al. 2024). However, considering the kinematic state of local dwarf galaxies in isolation, including gas-poor ones (Taibi et al. 2018, 2020), this possibility seems less likely (Leaman et al. 2017; Wheeler et al. 2017).
Our results on IC 1613 show the importance of performing a multi-tracer chemo-kinematic analysis covering several age ranges. Ongoing and future spectroscopic surveys, such as DESI (Levi et al. 2013), WEAVE (Jin et al. 2024) and 4MOST (de Jong et al. 2019), will provide such an analysis for the nearest dwarf galaxies with large statistical samples. At the distance of IC 1613 and beyond, we have shown here the challenges inherent in obtaining large samples covering a wide-area, for which we should await the arrival of planned or proposed facilities, such as the Mauna Kea Spectroscopic Explorer (MSE The MSE Science Team et al. 2019) and the Wide-field Spectroscopic Telescope (WST Mainieri et al. 2024).
6.2 Dynamical mass estimation and the role of inclination
We previously mentioned that the HI inclination value of IC 1613 is a matter of debate. Using low-resolution HI data, Lake & Skillman (1989) reported a value of , in very good agreement with the best-fitting value from Read et al. (2016, see Table 1), which we have adopted. A similar value was also obtained from Oh et al. (2015), although unconstrained. However, Read et al. (2016) suggested that the true inclination could be as low as , which allowed to bring the galaxy in agreement with model expectations. In fact, the HI circular velocity curve of IC 1613 grows very slowly, even compared to expectations of a cored dark matter density profile and assuming the most favourable value of low inclination. This appears to be the consequence of depressions and holes in the HI field, with the stellar field probably having been affected by them too, as we saw in Sects. 4.2 and 4.3. For Collins & Read (2022) an inclination value of would bring the galaxy in agreement with expectations from the baryonic Tully-Fisher relation (BTFR). We note that such a low inclination is at odds with Lake & Skillman (1989), who argue that a higher inclination value is to be preferred on the basis that the external isophotes of the HI (assumed as a thin disk) agree well in shape and orientation with those of the stellar component, as is also the case between the spatial and kinematic P.A. of the HI. Indeed from the stellar photometric axis ratio we would obtain an inclination of . If the galaxy was seen almost face-on, it would be difficult to explain such features.
Here we attempt to derive the inclination angle entirely from the stellar component, using the RGS and MS stars as independent dynamical tracers, whose properties have been reported in Sect. 4.3. The idea is quite simple: since both tracers must share the same circular velocity, we can equalise the formula for the asymmetric drift correction for both of them and solve for the inclination. In practice, we equalise the right-hand side of the following equation: , where is the asymmetric drift correction term necessary to remove the random motion component that suppresses the rotation curve. Another implicit assumption is that both tracers also share the same inclination angle. The term could be rather complex, as it takes into account the shape of the velocity ellipsoid and the tracer density profile (see e.g. Weijmans et al. 2008).
As a first attempt, we follow the formalism of Read et al. (2016) which primarily assumes a kinematic tracer with an exponential surface brightness profile and a constant observed velocity dispersion profile. This reduces the above equation to: , where is the exponential scale radius. We further assume that both tracers share the same scale radius (as ). To solve for the inclination, we draw random values from the distributions of the rotation and velocity dispersion values, assuming Gaussian errors. We find then an inclination of . Such value is dominated by the observational errors, in particular of the MS stars.
We also attempted a more sophisticated approach following, for instance, the prescription of Weijmans et al. (2008, using their Eqs. 6-12 and following considerations in their Appendix A2). In particular, assuming that the galaxy is axisymmetric, that the rotation curve flattens at large radii and that the potential is spherical, the asymmetric drift correction term becomes:
where the first term between parenthesis is the slope of the surface brightness profile, again assumed to be exponential, while the last two terms are connected to the shape of the velocity ellipsoid and the shape of the rotation curve. The former is modelled with a parameter , while the latter assumes that the rotation curve follows a power-law model where the curve becomes flat at large radii. The radius at which the rotation curve starts to flatten is called core radius and enters in the equation as . We calculate from the HI rotation curve ( arcmin) and assume it is the same also for the stellar tracers. While calculating , we also randomise over these two terms assuming a uniform distribution for and a Gaussian one for . Nevertheless, their contribution to the final value resulted to be minor. The radial dispersion is instead parametrised as , which can be obtained from the observed following Eq. (11) of Weijmans et al. (2008). Such equation depends on , but also on and . Therefore, in order to find a parametric solution of , we also randomise over that will act as a prior distribution, which we assume uniform over , with . Compared to the simpler approach described above, we find here a slightly lower inclination value, , but a comparable uncertainty.
Both recovered inclinations lie in the middle of the literature values, with the error bars preventing one from favouring a high (Lake & Skillman 1989) or low value (Read et al. 2016; Collins & Read 2022). It seems though that a value as low as is less favoured. If an inclination of will be preferred, this will confirm IC 1613 as an outlier of the BTFR. Such behaviour seems unusual for a galaxy with a baryonic mass of (Read et al. 2016), while it seems to resemble that of gas-rich ultra-diffuse galaxies found in the field (though an order of magnitude more massive; see Mancera Piña et al. 2019). We defer to a future dynamical study the task of better quantifying the value of inclination, where all available kinematic tracers, be they gas and stars, are simultaneously taken into account (see e.g. Leung et al. 2021, for WLM).
We can now estimate IC 1613’s dynamical mass, based on the stellar rotational velocities, corrected for the galaxy’s inclination, and the asymmetric drift; therefore turned into circular velocities . For the inclination, we stick to the last calculated value of , while for simplicity we assume again the asymmetric drift correction from Read et al. (2016) in order to estimate . We perform this estimate within the 3D de-projected half-light radius to make a direct comparison with values from mass estimators for dispersion supported systems. We take the average of the observed values of and between F2 and F3, assuming that both the rotation and dispersion profiles are slowly varying up to the half-light radius. The circular velocity is then km s-1. The dynamical mass is then , while the corresponding mass-to-light ratio within the half-light radius would be .
For comparison, the Wolf et al. (2010) mass-estimator for dispersion-supported systems, that is , where G is the gravitational constant, yields within the same radius. This is within the previous results accounting for inclination, but also within the value reported by Kirby et al. (2014). We note that this mass estimator is valid as long as the system is spherically symmetric, without rotation and with a flat l.o.s. velocity dispersion profile. To some degree, all these conditions are violated for IC 1613.
@@ 7 Summary and future work
The quality of the dataset allowed for a generally accurate spectral classification (i.e., with determined to better than 500 K) of 808 stars. We found a majority of K-type stars, and representatives from all the other types (from O-type to M-type stars). The classification included the identification of a sample of 24 probable Be stars, together with a sample of 14 probable C stars.
@@
For all types except the Be stars, the l.o.s. velocities have been determined with a pixel-by-pixel analysis, resulting in measurements with a mean precision of km s-1 and signal-to-noise ratio S/N. For the Be stars we have determined the velocities by cross-correlation with dedicated templates. From this sample we identified 746 probable members (i.e., having ), which show a significant rotation pattern. When modelled as a linear velocity gradient this yields km s-1 arcmin-1, along the HI kinematic major axis (with P. A. ). This is the first time that rotation for the stellar component is detected with high significance for this galaxy.
@@
We further analysed the probable members looking at the kinematic properties of each individual field and as a function of stellar types tracing different broad age ranges. The individual fields analysis confirmed the presence of a rotation signal, and a velocity dispersion profile that decreases with radius. On average, the stellar component shows higher velocity dispersion and a slower rotation than the HI gas; this is a manifestation of the asymmetric drift, as demonstrated by the fact that the kinematics of the two tracers comes in good agreement when applying an asymmetric drift correction. The kinematic analysis of the different selected stellar populations showed a coherent picture with the analysis of the main sample, with the MS stars following more closely the kinematics of the HI gas than the evolved red giant stars.
@@
The chemical analysis was conducted on the sub-sample of RGB stars exploiting the wavelength region of the Ca II triplet lines. We obtained a median dex, with scatter values of dex and dex, in good agreement with previous results reported by Kirby et al. (2013). The average [Fe/H] is well within the rms scatter of the stellar luminosity-metallicity relation found for the LG dwarf galaxies by the same authors. The spatial distribution of the [Fe/H] values showed no significant signs of a radial metallicity gradient, compatible with results from other LG dwarf galaxies (Leaman et al. 2013; Taibi et al. 2022).
Examining the kinematics of stars as a function of broad age ranges, we find that the velocity dispersion increases as a function of age, with the behaviour being very clear in the outermost pointings, while the rotation-to-velocity dispersion support decreases. On timescales of Gyr, the stellar kinematics still follow very closely that of the neutral gas, while the two components decouple on longer timescales. We interpret these results as a consequence of stars being born from a less turbulent gas over time. On the other hand, with the dynamical mass estimation we provide a new value of the inclination angle of IC 1613. To this end, we use stellar tracers with different observed kinematics that must be consistent with the same underlying gravitational potential.
@@
In future efforts, the data presented here could be used to perform a dedicated analysis of the massive stars in our sample using the FASTWIND stellar atmosphere code (Puls et al. 2005), which will allow us to obtain reliable stellar parameters from their spectra. We also aim to improve the identification process of the M-type stars in our sample, and therefore to constrain the spectroscopic C/M ratio. This will tell us more of the chemical properties of the intermediate-age stellar population of IC 1613. This MUSE data-set could also be used to analyse the ionised gas medium using available emission lines, like or nebular oxygen lines in order to get an insight on the kinematics of the shell-like gas structure visible in field F1 and compare it with that of the youngest stars. Finally, we aim to perform a dedicated dynamical study where all available kinematic tracers, be they gas and stars, are simultaneously taken into account in order to firmly establish the (dark) matter distribution of IC 1613.
@@ Acknowledgements.
We wish to thank the anonymous referee for the constructive comments that helped to improve the manuscript. We also thank M.-R. Cioni, L. R. Patrick, J. Read, F. Lelli, and I. Trujillo for useful discussions and comments at various stage of this project. ST acknowledges funding of a Leibniz-Junior Research Group (PI: M. Pawlowski; project number J94/2020) via the Leibniz Competition. GB and CG acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia en Innovación (AEI-MICIN) and the European Regional Development Fund (ERDF) under Grant Number PID2020-118778GB-I00/10.13039/50110001103 and the AEI under Grant Number CEX2019-000920-S. GI acknowledges financial support from the European Research Council (ERC) through the ERC Consolidator Grant DEMOBLACK, under contract No. 770017. SK acknowledges funding from UKRI in the form of a Future Leaders Fellowship (grant no. MR/T022868/1). PEMP acknowledges the support from the Dutch Research Council (NWO) through the Veni grant VI.Veni.222.364. This research has made use of NASA’s Astrophysics Data System, VizieR catalogue access tool (CDS, Strasbourg, France, DOI: 10.26093/cds/vizier), and extensive use of Python3.8 (Van Rossum & Drake 2009), including iPython (v8.12, Pérez & Granger 2007), Numpy (v1.24, Harris et al. 2020), Scipy (v1.10, Virtanen et al. 2020), Matplotlib (v3.7, Hunter 2007), Astropy (v5.2, Astropy Collaboration et al. 2022) and Scikit-learn (v1.2, Pedregosa et al. 2011) packages.References
@@
- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
- de Jong et al. (2019) de Jong, R. S., Agertz, O., Berbel, A. A., et al. 2019, The Messenger, 175, 3
- de los Reyes et al. (2023) de los Reyes, M. A. C., Kirby, E. N., Zhuang, Z., et al. 2023, ApJ, 951, 52
@@
- Dorman et al. (2015) Dorman, C. E., Guhathakurta, P., Seth, A. C., et al. 2015, ApJ, 803, 24
@@
- Forbes et al. (2012) Forbes, J., Krumholz, M., & Burkert, A. 2012, ApJ, 754, 48
- Fraternali et al. (2021) Fraternali, F., Karim, A., Magnelli, B., et al. 2021, A&A, 647, A194
@@
- Giesers et al. (2018) Giesers, B., Dreizler, S., Husser, T.-O., et al. 2018, MNRAS, 475, L15
- Giesers et al. (2019) Giesers, B., Kamann, S., Dreizler, S., et al. 2019, A&A, 632, A3
@@
- Lelli et al. (2021) Lelli, F., Di Teodoro, E. M., Fraternali, F., et al. 2021, Science, 371, 713
- Leung et al. (2021) Leung, G. Y. C., Leaman, R., Battaglia, G., et al. 2021, MNRAS, 500, 410
- Levi et al. (2013) Levi, M., Bebek, C., Beers, T., et al. 2013, arXiv e-prints, arXiv:1308.0847
@@
- Mainieri et al. (2024) Mainieri, V., Anderson, R. I., Brinchmann, J., et al. 2024, arXiv e-prints, arXiv:2403.05398
- Mancera Piña et al. (2019) Mancera Piña, P. E., Fraternali, F., Adams, E. A. K., et al. 2019, ApJ, 883, L33
- Mancera Piña et al. (2024) Mancera Piña, P. E., Golini, G., Trujillo, I., & Montes, M. 2024, arXiv e-prints, arXiv:2404.06537
- Mancera Piña et al. (2021) Mancera Piña, P. E., Posti, L., Fraternali, F., Adams, E. A. K., & Oosterloo, T. 2021, A&A, 647, A76
@@
- Pessa et al. (2023) Pessa, I., Schinnerer, E., Sanchez-Blazquez, P., et al. 2023, A&A, 673, A147
- Poci et al. (2019) Poci, A., McDermid, R. M., Zhu, L., & van de Ven, G. 2019, MNRAS, 487, 3776
@@
- Quirk et al. (2019) Quirk, A., Guhathakurta, P., Chemin, L., et al. 2019, ApJ, 871, 11
- Quirk et al. (2022) Quirk, A. C. N., Guhathakurta, P., Gilbert, K. M., et al. 2022, AJ, 163, 166
@@
- Rizzo et al. (2021) Rizzo, F., Vegetti, S., Fraternali, F., Stacey, H. R., & Powell, D. 2021, MNRAS, 507, 3952
- Rizzo et al. (2020) Rizzo, F., Vegetti, S., Powell, D., et al. 2020, Nature, 584, 201
@@
- Spitzer & Schwarzschild (1951) Spitzer, Lyman, J. & Schwarzschild, M. 1951, ApJ, 114, 385
- Spitzer & Schwarzschild (1953) Spitzer, Lyman, J. & Schwarzschild, M. 1953, ApJ, 118, 106
@@
- Taibi et al. (2018) Taibi, S., Battaglia, G., Kacharov, N., et al. 2018, A&A, 618, A122
@@
- The MSE Science Team et al. (2019) The MSE Science Team, Babusiaux, C., Bergemann, M., et al. 2019, arXiv e-prints, arXiv:1904.04907
@@
- Wisnioski et al. (2019) Wisnioski, E., Förster Schreiber, N. M., Fossati, M., et al. 2019, ApJ, 886, 124
- Wisnioski et al. (2015) Wisnioski, E., Förster Schreiber, N. M., Wuyts, S., et al. 2015, ApJ, 799, 209